an investor invested a total of 1,900.00 in two mutual funds. One fund earned a 9% profit while the other earned a 2% profit. If the investor's total profit was 52.00, how much was invested in each mutual fund?

thank you in advance!!!!!!!!! im so illiterate in math

Answer by jim_thompson5910(35256)

Let x = amount invested that generated 9% profit and y = amount invested that generated 2% profit

Since "an investor invested a total of 1,900.00 in two mutual funds", we know that

x+y = 1900

and because "One fund earned a 9% profit while the other earned a 2% profit. If the investor's total profit was 52.00", we know that

0.09x + 0.02y = 52

-------------------------------------------------------

Let's use these equations to solve for x and y

First solve the first equation for y

x+y = 1900

y = 1900 - x

Now plug this into the second equation

0.09x + 0.02y = 52

0.09x + 0.02(1900 - x) = 52

0.09x + 0.02(1900) + 0.02(-x) = 52

0.09x + 38 - 0.02x = 52

0.07x + 38 = 52

0.07x = 52 - 38

0.07x = 14

x = 14/(0.07)

x = 200

Now use this to find y

y = 1900 - x

y = 1900 - 200

y = 1700

So x = 200 and y = 1700, which means that $200 generated a 9% profit and the other $1700 generated a 2% profit.

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------